# Neural Force Field: Few-shot Learning of Generalized Physical Reasoning

### v4

@@ 4.1 Environments and datasets
Figure 7: Trajectory predictions on unseen scenarios after few-shot learning. After learning from 100 and 200 trajectories on I-PHYRE and N-body systems, respectively, our NFF predictions closely match the ground truth behaviors across diverse scenarios, from rigid body interactions to gravitational dynamics. In contrast, other baselines fail to predict physically plausible dynamics. Additional visualizations are provided in Appx.˜G.
@@ 4.2 Learning force fields from a few examples
We qualitatively evaluate NFF’s ability to inverse force fields from limited observations without direct access to ground-truth forces; training details are provided in Appx.˜B. For I-PHYRE, we train NFF on 10 basic games with 10 trajectory samples each. From these 100 trajectories, the model successfully learns fundamental physical concepts through unified force fields. To verify that NFF learns generalizable physical principles, we examine force responses to varied ball-block interactions under controlled conditions. Fig.˜A1 shows our systematic evaluation where we independently vary individual parameters such as position, velocity, and angle while holding others constant, demonstrating NFF’s accurate force response modeling. For the N-body system, we train NFF using 200 randomly sampled trajectories from 2-body and 3-body dynamics. As shown in Fig.˜5(b), the model successfully learns to capture the inverse gravitational field, correctly modeling the distance-dependent centripetal forces governing the mutual attraction between bodies. For PHYRE, we train a vision-based NFF on 12,00012{,}000 trajectories (20 templates $\times$ 20 tasks $\times$ 30 actions), which is about 267 times less data than prior SOTA (Qi et al., , 2021). As shown in Fig.˜5(a), the model accurately captures force fields across diverse templates.
Table 1: NFF outperforms baselines on prediction tasks. Comparison of NFF, IN (Battaglia et al., , 2016), GCN (Kipf and Welling, , 2016), and SlotFormer (Wu et al., , 2023) on I-PHYRE and N-body tasks. Results are reported with SEM; lower is better ($\downarrow$) for error metrics (RMSE, FPE, PCE), and higher is better ($\uparrow$) for correlation (R). NFF consistently yields lower errors and higher correlations, especially under cross-scenario generalization. The [0,3T] settings extend prediction from 50 to 150 steps.
@@
RMSE $\downarrow$ IN 0.124 ± 0.007 0.194 ± 0.004 0.306 ± 0.022 0.832 ± 0.034 4.727 ± 0.111
GCN 0.118 ± 0.008 0.213 ± 0.006 0.387 ± 0.028 1.090 ± 0.050 2.130 ± 0.046
Slotformer 0.067 ± 0.006 0.206 ± 0.005 0.257 ± 0.021 1.117 ± 0.042 3.531 ± 0.128
NFF 0.048 ± 0.004 0.131 ± 0.004 0.139 ± 0.012 0.558 ± 0.024 1.090 ± 0.023
FPE $\downarrow$ IN 0.129 ± 0.010 0.178 ± 0.005 0.283 ± 0.019 0.740 ± 0.026 5.376 ± 0.152
GCN 0.118 ± 0.009 0.213 ± 0.006 0.332 ± 0.023 1.028 ± 0.048 2.147 ± 0.046
Slotformer 0.068 ± 0.007 0.198 ± 0.005 0.250 ± 0.018 1.289 ± 0.066 3.708 ± 0.134
NFF 0.042 ± 0.004 0.111 ± 0.004 0.147 ± 0.013 0.506 ± 0.022 1.166 ± 0.024
PCE $\downarrow$ IN 0.004 ± 0.000 0.005 ± 0.000 0.018 ± 0.002 0.034 ± 0.003 0.060 ± 0.002
GCN 0.003 ± 0.000 0.194 ± 0.000 0.021 ± 0.002 0.034 ± 0.002 0.041 ± 0.002
Slotformer 0.002 ± 0.000 0.004 ± 0.000 0.015 ± 0.002 0.039 ± 0.002 0.075 ± 0.002
NFF 0.001 ± 0.000 0.003 ± 0.000 0.008 ± 0.001 0.022 ± 0.002 0.032 ± 0.002
R $\uparrow$ IN 0.993 ± 0.001 0.974 ± 0.001 0.968 ± 0.004 0.870 ± 0.011 0.325 ± 0.020
GCN 0.991 ± 0.001 0.969 ± 0.002 0.956 ± 0.005 0.834 ± 0.013 0.586 ± 0.023
Slotformer 0.997 ± 0.001 0.967 ± 0.001 0.977 ± 0.003 0.870 ± 0.010 0.420 ± 0.024
NFF 0.997 ± 0.001 0.980 ± 0.002 0.992 ± 0.001 0.938 ± 0.007 0.826 ± 0.011
@@ 4.3 Prediction on unseen scenarios
We evaluate the learned force fields’ predictive accuracy in both within- and cross-scenario settings. For I-PHYRE, we test against 20 ground truth trajectories per game, while for N-body, we evaluate using 200 unseen initial conditions. We compare against established interaction modeling methods: IN (Battaglia et al., , 2016), GCN (Kipf and Welling, , 2016), and SlotFormer (Wu et al., , 2023).
Evaluation metrics  We employ multiple complementary metrics to assess prediction quality. Beyond standard Root Mean Squared Error (RMSE), we introduce Final Position Error (FPE) and Position Change Error (PCE) for detailed performance analysis. FPE quantifies terminal position accuracy, while PCE evaluates the model’s ability to capture motion dynamics. Additionally, Pearson Correlation Coefficient (R) measures trajectory shape alignment independent of speed variations. This comprehensive metric suite enables thorough evaluation across temporal and spatial dimensions. Detailed definitions are provided in Sec.˜B.2.
Results  As shown in Fig.˜7, NFF generates physically plausible trajectories that closely match ground truth behavior, even in previously unseen scenarios. Quantitative results in Tab.˜1 demonstrate NFF’s superior performance across all metrics for both I-PHYRE and N-body tasks, with particularly strong results in cross-scenario generalization. While other models exhibits overfitting tendencies that limit the cross-scenario performance, NFF’s ability to learn generalizable physical principles enables robust prediction across diverse scenarios. Additional results in Appx.˜D demonstrate the superior robustness of our model under noisy state conditions.
@@ 4.4 Planning on unseen scenarios
We quantitatively evaluate planning by calculating success probability $p_{i}$ for each game as the success rate over 20 trials in round $i$, with NFF updating after failures; see detailed results in Appx.˜F. Figure˜10 compares NFF against random sampling, human performance (from Li et al., (2023)), IN, and SlotFormer using cumulative success probability: $1-\prod_{i=1}^{n}(1-p_{i})$ for current round $n$. NFF outperforms baselines and approaches human-level performance after refinement (Allen et al., , 2020), demonstrating effective few-shot learning. The performance of IN and SlotFormer, falling below random sampling, indicates how inaccurate dynamics modeling compromises planning.
Figure 10: Interactive planning performance on I-PHYRE. Comparison of cumulative success probability over 5 planning rounds among human, random sampling, our NFF, IN, and SlotFormer. Error bars show SEM across trials. Our NFF with refinement (Allen et al., , 2020) shows continuous improvement across rounds, achieving performance comparable to human (data from Li et al., (2023)), while others show limited performance even with refinement.
N-body planning  We focus on determining initial conditions that achieve desired final configurations under celestial dynamics. NFF’s inverse simulation enables direct computation of initial conditions through reverse time evolution, while IN and SlotFormer rely on iterative gradient-based refinement. As shown in Tab.˜3, NFF achieves superior planning performance in both within- and cross-scenario settings.
PHYRE planning  In PHYRE, we extend NFF to learn dynamics directly from RGB videos with segmentation masks by applying sparse convolutions to extract geometric features from the masked objects. Following PHYRE’s standard evaluation protocol, we first utilize NFF to predict trajectories and then use a pretrained TimeSFormer (Bertasius et al., , 2021) as the classifier to determine task success or failure. Planning performance is measured by the AUCCESS metric (Bakhtin et al., , 2019), which captures both effectiveness and efficiency by assigning higher weight to solutions that succeed with fewer attempts. Specifically, each attempt count $k\in{1,2,3,\dots,100}$ is weighted by $\omega_{k}=\log(k+1)-\log(k)$, and the final AUCCESS score is computed as: $\frac{\sum_{k}\omega_{k}\cdot s_{k}}{\sum_{k}\omega_{k}}$, where $s_{k}$ is the success rate within $k$ attempts.
We compare NFF against the RPIN (Qi et al., , 2021) and SlotFormer (Wu et al., , 2023). As shown qualitatively in Fig.˜9, NFF, trained on just 0.012 million trajectories, learns physically grounded dynamics that support effective planning while maintaining higher fidelity and stronger object permanence than RPIN and SlotFormer, particularly in cross-scenario generalization. Quantitative results in Tab.˜3 further demonstrate that NFF outperforms state-of-the-art baselines in the challenging cross-scenario tasks. In contrast, SlotFormer’s slot attention leads to limited generalization in unseen objects, resulting in planning failures in cross scenarios; see Fig.˜A8 for more analysis.
@@ 4.5 Ablation study
We investigate three key factors that may affect NFF performance, including integration precision, ODE grounding, and NOL. For integration precision, we evaluate NFF on the N-body task using Euler integration at four precision levels: 1e31e-3, 5e35e-3, 1e21e-2, and 5e25e-2. Tab.˜4 demonstrates that higher integration precision consistently yields better performance. We also explore more integration methods in Appx.˜C. To study ODE grounding’s impact on generalization, we compare two configurations: our NFF using predicted force fields with ODE integration, and IN using learned state transitions at matching step sizes. Tab.˜4 shows that ODE integration enhances generalization performance. For NOL, we replace the DeepONet in NFF with a vanilla MLP. The results in the last column of Tab.˜4 show a consistent increase in generalization errors across settings compared with NFF with DeepONet (column one). These ablation studies demonstrate that NFF achieves optimal few-shot learning of physical dynamics when combining high integration precision, ODE-based integration, and NOL.
Table 4: Ablation results emphasize the need for high integration precision, ODE grounding, and NOL. We test NFF with different integration precision. The NFF without ODE grounding degenerates to IN with the same integration precision (1e-3) and the NFF without NOL utilize an MLP as force predictor instead of DeepONet. NFF with high integration precision, ODE, and NOL yields the best generalization results.
Scenario NFF 1e-3 NFF 5e-3 NFF 1e-2 NFF 5e-2 w/o ODE w/o NOL
Training 0.083±0.009 0.137±0.014 0.129±0.016 0.231±0.022 0.044±0.005 0.056±0.005
Within 0.095±0.008 0.139±0.013 0.136±0.014 0.239±0.020 0.354±0.023 0.120±0.021
Within [0,3T] 0.471±0.025 0.558±0.024 0.535±0.027 0.693±0.030 0.949±0.041 0.540±0.042
Cross 1.024±0.021 1.090±0.023 1.084±0.024 1.315±0.033 3.518±0.061 1.347±0.128
@@ 5 Discussion
Modeling diverse forces  While NFF has demonstrated effectiveness in modeling fundamental physical forces like gravity, support, and collision, its principles extend naturally to broader domains. The framework can potentially model social forces (Shu et al., , 2015; Xie et al., , 2017; Wei et al., , 2017, 2018; Shu et al., , 2021) and biological forces governing migration patterns and human motion (Netanyahu et al., , 2024). Limitations  While our experiments focus on generalization across varying object masses, training a single model with varying friction and elasticity may pose additional challenges. Furthermore, for a clear and constrained validation of learning forces, we assume a deterministic environment with rigid bodies and have not explored the model’s performance in stochastic environments or those containing soft bodies and fluids. Impact statements  The proposed NFF framework could enable more sample-efficient and interpretable physical modeling, reducing data needed for training. We acknowledge potential risks if misused for simulating harmful physical scenarios. However, the research focuses on basic physical principles and publicly available phenomena, prioritizing transparent and reproducible science.
@@ 6 Conclusion
We present NFF, a force field-based representation framework for modeling complex physical interactions that exhibits human-like few-shot learning, generalization, and reasoning capabilities. Our experiments on three challenging physical reasoning datasets demonstrate NFF’s ability to not only learn diverse physical concepts and rules from limited observations but also generalize to unseen scenarios for both prediction and planning tasks. This initial exploration of the force field may provide new perspectives for developing physical world models through representation learning, potentially bridging the gap between human physical understanding and learning-based methods.
Acknowledgments and Disclosure of Funding

### v5

@@ 4.1 Environments and datasets
Figure 7: Trajectory predictions on unseen scenarios after few-shot learning. After learning from 100 and 200 trajectories on I-PHYRE and N-body systems, respectively, our NFF predictions closely match the ground truth behaviors across diverse scenarios, from rigid body interactions to gravitational dynamics. In contrast, other baselines fail to predict physically plausible dynamics. Additional visualizations are provided in Appendix H.
@@ 4.2 Learning force fields from a few examples
We qualitatively evaluate NFF’s ability to inverse force fields from limited observations without direct access to ground-truth forces; training details are provided in Appendix C. For I-PHYRE, we train NFF on 10 basic games with 10 trajectory samples each. From these 100 trajectories, the model successfully learns fundamental physical concepts through unified force fields. To verify that NFF learns generalizable physical principles, we examine force responses to varied ball-block interactions under controlled conditions. Figure A1 shows our systematic evaluation where we independently vary individual parameters such as position, velocity, and angle while holding others constant, demonstrating NFF’s accurate force response modeling. For the N-body system, we train NFF using 200 randomly sampled trajectories from 2-body and 3-body dynamics. As shown in Figure 5(b), the model successfully learns to capture the inverse gravitational field, correctly modeling the distance-dependent centripetal forces governing the mutual attraction between bodies. For PHYRE, we train a vision-based NFF on 12,00012{,}000 trajectories (20 templates $\times$ 20 tasks $\times$ 30 actions), which is about 267 times less data than prior SOTA (Qi et al., 2021). As shown in Figure 5(a), the model accurately captures force fields across diverse templates.
Table 1: NFF outperforms baselines on prediction tasks. Comparison of baselines on I-PHYRE and N-body tasks. Results are reported with SEM; lower is better ($\downarrow$) for error metrics (RMSE, FPE, PCE), and higher is better ($\uparrow$) for correlation (R). NFF consistently yields lower errors and higher correlations, especially under cross-scenario generalization. The [0,3T] settings extend prediction from 50 to 150 steps.
@@
RMSE $\downarrow$ IN 0.124 ± 0.007 0.194 ± 0.004 0.200 ± 0.015 0.752 ± 0.026 6.942 ± 0.235
Slotformer 0.067 ± 0.006 0.206 ± 0.005 0.214 ± 0.018 1.092 ± 0.043 2.533 ± 0.068
SGENO 0.203 ± 0.070 0.314 ± 0.106 0.079 ± 0.009 0.690 ± 0.144 2.759 ± 0.076
NFF 0.048 ± 0.004 0.131 ± 0.004 0.097 ± 0.009 0.525 ± 0.026 1.226 ± 0.023
FPE $\downarrow$ IN 0.129 ± 0.010 0.178 ± 0.005 0.198 ± 0.014 0.785 ± 0.027 11 .568 ± 0.492
Slotformer 0.068 ± 0.007 0.198 ± 0.005 0.215 ± 0.017 1.311 ± 0.064 2.751 ± 0.060
SEGNO 0.212 ± 0.082 0.289 ± 0.109 0.074 ± 0.009 1.535 ± 0.056 2.302 ± 0.062
NFF 0.042 ± 0.004 0.111 ± 0.004 0.104 ± 0.009 0.592 ± 0.029 1.418 ± 0.029
PCE $\downarrow$ IN 0.004 ± 0.000 0.005 ± 0.000 0.012 ± 0.001 0.029 ± 0.002 0.101 ± 0.003
Slotformer 0.002 ± 0.000 0.004 ± 0.000 0.013 ± 0.002 0.035 ± 0.002 0.053 ± 0.002
SEGNO 0.008 ± 0.008 0.017 ± 0.018 0.005 ± 0.001 0.022 ± 0.004 0.063 ± 0.001
NFF 0.001 ± 0.000 0.003 ± 0.000 0.007 ± 0.001 0.021 ± 0.002 0.035 ± 0.002
R $\uparrow$ IN 0.993 ± 0.001 0.974 ± 0.001 0.984 ± 0.002 0.895 ± 0.009 0.287 ± 0.015
Slotformer 0.997 ± 0.001 0.967 ± 0.001 0.977 ± 0.003 0.854 ± 0.010 0.4384 ± 0.019
SEGNO 0.977 ± 0.026 0.928 ± 0.064 0.996 ± 0.001 0.947 ± 0.010 0.291 ± 0.013
NFF 0.997 ± 0.001 0.980 ± 0.002 0.996 ± 0.001 0.949 ± 0.006 0.803 ± 0.010
@@ 4.3 Prediction on unseen scenarios
We evaluate the learned force fields’ predictive accuracy in both within- and cross-scenario settings. For I-PHYRE, we test against 20 ground truth trajectories per game, while for N-body, we evaluate using 200 unseen initial conditions. We compare against established state-of-the-art (SOTA) interaction modeling methods: vanilla IN (Battaglia et al., 2016), SEGNO (Liu et al., 2024b), and SlotFormer (Wu et al., 2023). Note that for fair comparison, we directly feed encoded state information into SlotFormer to avoid inaccurate perception caused by slot attention. More baseline performances on N-body dataset can be found in Appendix D
Evaluation metrics   We employ multiple complementary metrics to assess prediction quality. Beyond standard Root Mean Squared Error (RMSE), we introduce Final Position Error (FPE) and Position Change Error (PCE) for detailed performance analysis. FPE quantifies terminal position accuracy, while PCE evaluates the model’s ability to capture motion dynamics. Additionally, Pearson Correlation Coefficient (R) measures trajectory shape alignment independent of speed variations. This comprehensive metric suite enables thorough evaluation across temporal and spatial dimensions. Detailed definitions are provided in Section C.2.
Results   As shown in Figure 7, NFF generates physically plausible trajectories that closely match ground truth behavior, even in previously unseen scenarios. Quantitative results in Table 1 demonstrate NFF’s superior performance across all metrics for both I-PHYRE and N-body tasks, with particularly strong results in cross-scenario generalization. While other models exhibit overfitting tendencies that limit the cross-scenario performance, NFF’s ability to learn generalizable physical principles enables robust prediction across diverse scenarios.
@@ 4.4 Planning on unseen scenarios
We quantitatively evaluate planning by calculating success probability $p_{i}$ for each game as the success rate over 20 trials in round $i$, with NFF updating after failures; see detailed results in Appendix G. Figure 10 compares NFF against random sampling, human performance (from Li et al. (2023)), IN, and SlotFormer using cumulative success probability: $1-\prod_{i=1}^{n}(1-p_{i})$ for current round $n$. NFF outperforms baselines and approaches human-level performance after refinement (Allen et al., 2020), demonstrating effective few-shot learning. The performance of IN and SlotFormer, falling below random sampling, indicates how inaccurate dynamics modeling compromises planning.
Figure 10: Interactive planning performance on I-PHYRE. Comparison of cumulative success probability over 5 planning rounds among human, random sampling, our NFF, IN, and SlotFormer. Error bars show SEM across trials. Our NFF with refinement (Allen et al., 2020) shows continuous improvement across rounds, achieving performance comparable to human (data from Li et al. (2023)), while others show limited performance even with refinement.
N-body planning   We focus on determining initial conditions that achieve desired final configurations under celestial dynamics. NFF’s inverse simulation enables direct computation of initial conditions through reverse time evolution, while IN and SlotFormer rely on iterative gradient-based refinement. As shown in Table 3, NFF achieves superior planning performance in both within- and cross-scenario settings. Moreover, its reliance on backward integration provides a substantial speed advantage, as further evidenced by the planning time evaluation in Table A3.
PHYRE planning   In PHYRE, we extend NFF to learn dynamics directly from RGB videos with segmentation masks by applying sparse convolutions to extract geometric features from the masked objects. Following PHYRE’s standard evaluation protocol, we first utilize NFF to predict trajectories and then use a pretrained TimeSFormer (Bertasius et al., 2021) as the classifier to determine task success or failure. Planning performance is measured by the AUCCESS metric (Bakhtin et al., 2019), which captures both effectiveness and efficiency by assigning higher weight to solutions that succeed with fewer attempts. Specifically, each attempt count $k\in{1,2,3,\dots,100}$ is weighted by $\omega_{k}=\log(k+1)-\log(k)$, and the final AUCCESS score is computed as: $\frac{\sum_{k}\omega_{k}\cdot s_{k}}{\sum_{k}\omega_{k}}$, where $s_{k}$ is the success rate within $k$ attempts.
@@
We compare NFF against the RPIN (Qi et al., 2021) and SlotFormer (Wu et al., 2023). As shown qualitatively in Figure 9, NFF, trained on just 0.012 million trajectories, learns physically grounded dynamics that support effective planning while maintaining higher fidelity and stronger object permanence than RPIN and SlotFormer, particularly in cross-scenario generalization. Quantitative results in Table 3 further demonstrate that NFF outperforms SOTA baselines in the challenging cross-scenario tasks. In contrast, SlotFormer’s slot attention leads to limited generalization in unseen objects, resulting in planning failures in cross scenarios; see Figure A8 for more analysis.
@@ 4.5 Ablation study
Table 4: Ablation results emphasize the need for high integration precision, ODE grounding, and NOL. We test NFF with different integration precision. The NFF without ODE grounding degenerates to IN with the same integration precision (5e-3) and the NFF without NOL utilize an MLP as force predictor instead of DeepONet. NFF with high integration precision, ODE, and NOL yields the best generalization results.
Scenario NFF 1e-3 NFF 5e-3 NFF Adaptive w/o ODE w/o NOL
Training 0.080±0.009 0.107±0.013 0.072±0.008 0.044±0.005 0.056±0.005
Within 0.097±0.009 0.120±0.011 0.092±0.007 0.354±0.023 0.120±0.021
Within [0,3T] 0.525±0.026 0.557±0.025 0.613±0.037 0.949±0.041 0.540±0.042
Cross 1.226±0.023 1.251±0.025 1.788±0.375 3.518±0.061 1.347±0.128
We investigate three key factors that may affect NFF performance, including integration precision, ODE grounding, and NOL. For integration precision, we evaluate NFF on the N-body task using Euler integration at two precision levels: 1e31e-3, 5e35e-3 and adaptive integration with tolerance level: 1e51e-5. Table 4 demonstrates that higher integration precision consistently yields better performance. We also explore more integration methods in Appendix E. To study ODE grounding’s impact on generalization, we compare two configurations: our NFF using predicted force fields with ODE integration, and IN using learned state transitions at matching step sizes. Table 4 shows that ODE integration enhances generalization performance. For NOL, we replace the DeepONet in NFF with a vanilla MLP. The results in the last column of Table 4 show a consistent increase in generalization errors across settings compared with NFF with DeepONet (column one). These demonstrate that NFF achieves optimal few-shot learning of physical dynamics when combining high integration precision, ODE-based integration, and NOL.
@@ 5 Discussion
Extending NFF to real stimuli   While NFF has demonstrated effectiveness in modeling fundamental physical forces like gravity, support, and collision on synthetic datasets, extending NFF to real-world environments remains unexplored. A promising avenue lies in harnessing the strong generalization capabilities of vision foundation models to translate complex, real-world visual inputs into object-centric representations compatible with NFF. Despite the challenges of perceptual noise and diverse physical interactions, we believe this could enable the framework to ground high-level visual perception in explicit force-based dynamics modeling, effectively bridging the gap between video prediction and human-level physical reasoning.
Limitations   While our experiments focus on generalization across varying object masses and object compositions in controllable abstract reasoning tasks, training a single model with varying friction and elasticity may pose additional challenges. Furthermore, for a clear and constrained validation of learning forces, we assume a deterministic environment with rigid bodies and have not explored the model’s performance in stochastic environments or those containing soft bodies and fluids.
Impact statements   The proposed NFF framework could enable more sample-efficient and interpretable physical modeling, reducing data needed for training. We acknowledge potential risks if misused for simulating harmful physical scenarios. However, the research focuses on basic physical principles and publicly available phenomena, prioritizing transparent and reproducible science.
@@ 6 Conclusion
We present NFF, a force field-based representation framework for modeling complex physical interactions that exhibits human-like few-shot learning, generalization, and reasoning capabilities. Our experiments on three challenging abstract reasoning datasets with complex multi-object interactions demonstrate NFF’s ability to not only learn diverse physical concepts and rules such as gravity, collision, friction, and attraction from limited observations but also generalize to unseen scenarios for both prediction and planning tasks. This initial exploration of the force field may provide new perspectives for developing physical world models through representation learning, potentially bridging the gap between human physical understanding and learning-based methods.
Reproducibility statement
To support reproducibility, we provide detailed documentation on the data and environment configurations in Appendix B, describe the model implementation in Appendix C, explain the evaluation metrics in Section C.2, and present the ablation studies and additional results in Appendix D, Appendix E, Appendix F, and Appendix G. Both the datasets and code will be made publicly available.
Acknowledgments